3.      You are given the following MPC’s, for each one calculate (show work) the MPS and the Multiplier

a. MPC = .80      B. MPC = .40               c. MPC = .67      D. MPC = .33

4.      You are given the following MPS’s, for each one calculate (show work) the MPC and the Multiplier

a. MPS = .50      B. MPS = .75               c. MPS = .30      D. MPS = .25

5.      If the multiplier was equal to 2.50, explain what impact the following changes would have on GDP. Show all work.

a.       Consumption increases by $100B                  B. Investment declines by $8B          c. Imports increase by $5B
Answer #1

3.      Given the following MPC’s, for each one calculate (show work) the MPS and the Multiplier

a. MPC = .80  

MPS=1-MPC

=1-0.8

=0.2

Multiplier=1/MPS

=1/0.2

=5

b.

MPC = .40

MPS=1-MPC

=1-0.4

=0.6

Multiplier=1/MPS

=1/0.6

=1.66

c.

MPC = .67

MPS=1-MPC

=1-0.67

=0.33

Multiplier=1/MPS

=1/0.33

=3.03

d.

MPC = .33

MPS=1-MPC

=1-0.33

=0.67

Multiplier=1/MPS

=1/0.67

=1.49

4.

a.

MPS=0.5

MPC=1-MPS

=1-0.5

=0.5

Multiplier=1/0.5

=2

b.

MPS=0.75

MPC=1-MPS

=1-0.75

=0.25

Multiplier=1/0.75

=1.33

c.

MPS=0.3

MPC=1-MPS

=1-0.3

=0.7

Multiplier=1/0.3

=3.33

d.

MPS=0.25

MPC=1-MPS

=1-0.25

=0.75

Multiplier=1/0.25

=4

5.

If Multiplier=2.50

a.

If consumption increase by $100 B, then GDP (Y) will increase by=multiplier* change in C

=2.50*100 B

=$250B

b.

If investment decrease by $8 B, then GDP (Y) will decrease by=multiplier* change in I

=2.50*(-$8 B)

=$20 B

C.

If import increase by $5 B, then GDP (Y) will decrease by=multiplier* change in import

=2.50*5 B

=$12.50B